OpenStudy (yourboydusty):

Which point lies on the line described by the equation below? y - 3 = 2(x + 5) I have already deduced some of it to y-3=2(x+5) now what do i do? A. (3, -3) B. (4, 21) C. (4, 18) D. (3, 17) E. (-2, 13) F. (-1, 9)

OpenStudy (saadsword):

y=2(x+5)-3=2x+10-3=2x+7 y=2x+7

OpenStudy (saadsword):

now plug in the x value for each choice and see which one equals into the y value

OpenStudy (saadsword):

Choice A: y=2(3)+7=7+7=14 since plugging in 3 did not give -3 choice A is not correct Do the same with all the other choices to see which one is correct

OpenStudy (yourboydusty):

thank you ! :)